# Finding the particularity of the active episode of SGR J1935+2154 during which FRB 20200428 occurred: Implication from a statistics of Fermi/GBM X-ray bursts

### v1

@@ Abstract
By using the Fermi/GBM data of the X-ray bursts (XRBs) of SGR J1935+2154, we investigate the temporal clustering of the bursts and the cumulative distribution of the waiting time and fluence/flux. It is found that the bursts occurring in the episode hosting FRB 20200428 have obviously shorter waiting times than those in the other episodes. The general statistical properties of the XRBs further indicate they could belong to a self-organized critical (SOC) system (e.g., starquakes), making them very similar to the earthquake phenomena. Then, according to a unified scaling law between the waiting time and energy of the earthquakes as well as their aftershocks, we implement an analogy analysis on the XRBs and find that the FRB episode owns more dependent burst events than the other episodes. It is indicated that the FRB emission could be produced by the interaction between different burst events, which could correspond to a collision between different seismic/Alfven waves or different explosion outflows. Such a situation could appear when the magnetar enters into a global intensive activity period.
@@ 1 Introduction
Generally, magnetars are a class of neutron stars with an ultrahigh magnetic field ($\sim 10^{14}-10^{15}$ G), which usually exhibit in the Galaxy as pulsars of a long spin period ($2\sim 12$ s) and a fast spin-down rate ($\sim 10^{-13}-10^{-11}$ s s-1, Kouveliotou et al., 1998; Olausen & Kaspi, 2014). The marked feature of Galactic magnetars is their X-ray or soft $\gamma$-ray burst emission (Thompson & Duncan, 1995), which are observed in more than 2/3 of the magnetar population (Olausen & Kaspi, 2014).
@@
On 28th April 2020, from the direction of a Galactic magnetar, SGR J1935+2154, the Canadian Hydrogen Intensity Mapping Experiment (CHIME) captured a fast radio burst (FRB) emission (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), which was precisely temporally associated with an X-ray burst (XRB; Li et al., 2021, 2022a; Ge et al., 2022). This magnetar had just entered an active state of XRBs about half a month before FRB 20200428 occurred (Li et al., 2021; Cai et al., 2022a, b; Mereghetti et al., 2020; Younes et al., 2020; Tavani et al., 2020; Veres et al., 2020; Ridnaia et al., 2020; Barthelmy et al., 2020; Lin et al., 2020a; Ridnaia et al., 2021; Kaneko et al., 2021), and remained active at other times as well (Israel et al., 2016; Lin et al., 2020b; Borghese et al., 2022; Xie et al., 2022, 2023; Rehan & Ibrahim, 2023). Meanwhile, the dispersion measure (DM) of FRB 20200428 ($\sim 332.7$ pc cm -3) is well consistent with the distance of SGR J1935+2154 in the range of $6.6-12.5$ kpc (Kothes et al., 2018; Zhou et al., 2020; Zhong et al., 2020). These discoveries indicated that magnetars could be sources of the mysterious FRB phenomena (Lorimer et al., 2007; Thornton et al., 2013).
@@
In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular, magnetars (Popov & Postnov, 2010; Kulkarni et al., 2014; Katz, 2016; Connor et al., 2016; Cordes & Wasserman, 2016; Lyutikov, 2017a), which was firstly favored by the discovery of the repeating FRB 20121102 as well as its persistent radio counterpart (Kashiyama & Murase, 2017; Metzger et al., 2017; Cao et al., 2017; Dai & Lu, 2017; Michilli et al., 2018). Then, the direct connection of FRB 20200428 with SGR J1935+2154 undoubtedly offered a smoking-gun evidence for one of the FRB/FRB-like origins, the magnetar origin, although FRB 20200428 is still distinctive due to its obviously low luminosity than its cosmological cousins (CHIME/FRB Collaboration et al., 2020; Bochenek et al., 2020).
@@
Generally, the amounts of the energy releases of the bursts and outbursts of Galactic magnetars, from both soft gamma-ray repeaters (SGRs) and anomalous X-ray pulsars (AXPs), indicate that XRBs probably correspond to sudden and local (not global) catastrophic events. Since the energy is considered to eventually come from magnetic fields, the XRBs are most likely to be caused by sudden reordering of the magnetic field structure, which either deforms the crust drastically or/and leads to a major reconfiguration of the magnetosphere. As a result, an optically thick $e^{\pm}$ fireball can be expected to form, radiating X-rays (Thompson & Duncan, 1995, 2001; Beloborodov, 2021; Lyutikov, 2022). In the framework of such a starquake model, regardless of the specific X-ray radiation mechanism, some scale independent behaviors can be expected to appear in the statistics of the energy and waiting times of XRBs, just as found from earthquakes (Cheng et al., 1996). Many systems such as starquakes and earthquakes, consisting of a large number of entities that interact with each other in a complex way, can exhibit nonlinear behavior, which can be described by a self-organized criticality (SOC) theory. The complex interactions in such systems can occasionally lead to instabilities with subsequent small events. The largest events, the subsequent events, and the myriads of smaller events can finally share the same statistical property (Sornette, 2004; Tsallis, 2009; Aschwanden, 2011).
@@
Despite a basic consensus that could appear on the XRB origin, the mechanism responsible for the FRB emission is still under great debate. In the literature, two types of models have been suggested frequently, including the curvature radiation in the magnetosphere (Lu & Kumar, 2018; Yang & Zhang, 2018, 2020; Long & Pe’er, 2018; Wang et al., 2019; Kumar & Lu, 2020) and the shock synchrotron maser radiation (Lyubarsky, 2014; Beloborodov, 2017; Ghisellini, 2017; Kumar et al., 2017; Lyutikov, 2017b; Waxman, 2017; Metzger et al., 2019; Plotnikov & Sironi, 2019; Yu et al., 2021). In the former case, the curvature radiation could happen in the magnetosphere because the crust fracturing forms a charge-starved region and leads to the plasma discharges. In the latter case, the shock responsible for the synchrotron maser can be driven by the collisions between different ejecta of different burst activities. In observation, although FRB 20200428 was found to be in temporal coincidence with an XRB, the deep observations implemented by the Five Hundred Meter Aperture Telescope (FAST) on some other XRBs of SGR J1935+2154 ruled out the possibility of their association with FRB emission (Lin et al., 2020c). On the one hand, this fact could just be due to the narrow beaming of FRB emission, while the XRB emission is generally isotropic. In this case, there is in principle no intrinsic difference between the different XRBs. On the other hand, reversely, the un-association with FRB emission of other XRBs may indicate that the XRB associated with FRB 20200428 is intrinsically different from the others.
@@
Figure 1: The XRBs (triangle markers) of SGR J1935+2154 detected by GBM within different observation windows (horizontal lines). The length of the lines represents the duration of the windows. The first burst in each window is labelled by a black triangle, which is dropped in our statistics since its waiting time cannot be defined. Then, the windows that only have one burst are excluded (i.e, the grey lines). The remaining windows having sufficiently large number of bursts can be basically separated into four active episodes, as labelled by different colors (see Section 2.2 for active episode definition).
Figure 2: The corner plot of the shape parameter $k$ and the event rate $\lambda$ for the Weibull distribution (see Section 2.1).
@@
2 Temporal and energy distributions
We retrieve the burst start time, flux and fluence of SGR J1935+2154 which were detected by the Fermi/Gamma-ray Burst Monitor (GBM) from the published burst history (2014-2021, Table 4 in Lin et al. (2020b), Table 1 in Lin et al. (2020a), Table 1 in Kaneko et al. (2021), Table 1 in Rehan & Ibrahim (2023)). Then, we analyze the position history of Fermi/GBM observations using POSHIST  [^1] https://heasarc.gsfc.nasa.gov/FTP/fermi/data/gbm/daily/ and calculate the uninterrupted visible time (referred to as observation window hereafter) of the GBM to the SGR J1935+2154, excluding the Earth occultation and SAA (South Atlantic Anomaly) area, using the GBM Data Tools (Goldstein et al., 2022). The start times of the observation windows and hundreds of bursts are listed in Table 1 and shown in Fig. 1. To be specific, the burst samples of SGR J1935+2154 consist of 158 observation windows, with each window having an approximate duration of an hour because GBM orbits the Earth every about 90 minutes. However some observation windows are noticeably shorter than the 90 minutes because GBM occasionally shuts down when the satellite crosses the SAA area.
2.1 The temporal clustering of XRBs
@@
where $t_{i}$ is the burst start time of the $i-$th burst. Due to the unknown waiting time for the first burst in each observation window, only 171 out of a total of 329 bursts could be analyzed (see Fig. 1). A one-sample Kolmogorov-Smirnov test is performed to examine whether the waiting times follow an exponential distribution. The test yields a distance of $D=0.28$ and $p=0.03$, indicating that a non-homogeneous model is necessary to describe the temporal behavior of the bursts. Therefore, we use the Weibull function to model the waiting times of the non-stationary Poisson process. Specifically, for an observation window of a duration $T$ which owes a number of $N$ bursts, the likelihood of this window can be calculated by (Oppermann et al., 2018),
@@
×$\times\prod_{i=1}^{N-1}\mathcal{W}(t_{\mathrm{i}+1}-t_{\mathrm{i}}|k,\lambda)$
@@
Figure 3: The logarithm waiting times, $\log(\Delta{t})$, of the XRBs v.s their occurring times, where the time of FRB 20200428 is represented by the green dashed line. The number distribution of the bursts on the waiting times are exhibited in the right panel with the histograms, while the smooth lines give the kernel density of these distributions. The histogram distributions can be approximately fitted by a Gaussian function and its mean value and variance are represented by the orange dot and the short black line in the left panel.
2.2 The cumulative distribution of the waiting times
In Fig. 3, we plot the waiting times of the XRBs of SGR J1935+2154 v.s their burst times, where an apparent clustering of bursts appear. Therefore, we can divide these bursts into 4 separated active episodes as labelled by different colors, except several bursts (grey data) outside of these episodes. Here, an active episode can be defined as a period during which more than two bursts are produced within 10 days, as previously adopted in Lin et al. (2020b, a). The starting and ending times of the four active episodes are listed in Table 2. By according to the shape of the kernel of the histogram distributions of $\log_{10}(\Delta{t})$ shown in the right panel of Fig. 3, we tentatively use a Gaussian function to fit these histogram distributions. The obtained values of the mean and variance of the Gaussians are listed in Table 2. It can be found that the mean value of the waiting times in the 3rd episode (Apr.-May 2020) during which FRB 200428 occurred is significantly shorter than those in the other episodes, indicating the particularity of the FRB episode. In addition, these near-Gaussian distributions of the waiting times make these burst activities of SGR J1935+2154 similar to the previous burst activities of magnetars (Cheng et al., 1996; GöǧüŞ et al., 1999; Göǧüş et al., 2000; Younes et al., 2020; Cai et al., 2022a).
@@ 2.2 The cumulative distribution of the waiting times
For a more detailed description for the distribution property of the waiting times, we plot their complementary cumulative distribution (CCD) [^2] The cumulative distribution (CD) and CCD of the variable $x$ are defined in this paper as follows: $CD(x)$ $=P(\leq x)=\frac{n(\leq x)}{N}$ (5) $CCD(x)$ $=P(>x)=1-CD(x)$ where $N$ is the total number of the events. , $P(>\Delta{t})$, in the left panel of Fig. 4, where the distributions for different flux ranges are also displayed for comparison. According to the heavy-tail feature of these CCDs and following Reed & Jorgensen (2004), we try to use the double Pareto-Lognormal (dPLn) distribution to describe their general properties. The dPLn distribution, which consists of two Pareto tail distributions with a transition point in between, arises from a geometric Browning motion process, that will be killed when the killing rate comes to constant and the state becomes equilibrium (Mitzenmacher, 2004). Specifically, the dPLn description for the CCD of the waiting times can be expressed as (Reed & Jorgensen, 2004)
@@
where $\Phi$/$\Phi^{c}$ is the CD/CCD of standard normal distribution, $\nu$ and $\tau^{2}$ represent the mean and variance of the log-normal for the initial geometric Browning motion process, and
@@
The parameters $\alpha_{1}$ and $\alpha_{2}$ determine the slope of the two Pareto tails. By using the package distribuisonsrd of Comprehensive R Archive Network (CRAN, Mora-Cantallops et al., 2020), we perform the dPLn fitting to the CCDs of the waiting times, as shown by the dashed line in Figure 4, and obtain $\alpha_{1}$=3.85 and $\alpha_{2}$=0.73. This indicates that, in the range of a relatively long waiting time, the CCD can be approximated by a power law with a slope of $\alpha_{2}\sim 0.73$, which could further be independent of the selection of the flux threshold and thus of the burst energy.
@@
2.3 The cumulative distribution of the fluence/flux

### v6

# Finding the Particularity of the Active Episode of SGR J1935+2154 during Which FRB 20200428 Occurred: Implication from Statistics of Fermi/GBM X-Ray Bursts

@@ Abstract
By using the Fermi/Gamma-ray Burst Monitor data of the X-ray bursts (XRBs) of SGR J1935+2154, we investigate the temporal clustering of the bursts and the cumulative distribution of the waiting time and fluence/flux. It is found that the bursts occurring in the episode hosting FRB 20200428 have obviously shorter waiting times than those in the other episodes. The general statistical properties of the XRBs further indicate they could belong to a self-organized critical (SOC) system (e.g., starquakes), making them very similar to the earthquake phenomena. Then, according to a unified scaling law between the waiting time and energy of the earthquakes as well as their aftershocks, we implement an analogy analysis on the XRBs and find that the FRB episode owns more dependent burst events than the other episodes. It is indicated that the fast radio burst (FRB) emission could be produced by the interaction between different burst events, which could correspond to a collision between different seismic/Alfven waves or different explosion outflows. Such a situation could appear when the magnetar enters into a global intensive activity period.
@@ I Introduction
Generally, magnetars are a class of neutron stars with an ultrahigh magnetic field ($\sim 10^{14}-10^{15}$ G), which usually exhibit in the Galaxy as pulsars of a long spin period ($2\sim 12$ s) and a fast spin-down rate ($\sim 10^{-13}-10^{-11}$ s s-1, Kouveliotou1998; Olausen2014). The marked feature of Galactic magnetars is their X-ray or soft $\gamma$-ray burst emission (ThompsonDuncan1995), which are observed in more than 2/3 of the magnetar population (Olausen2014).
@@
On 28th April 2020, from the direction of a Galactic magnetar, SGR J1935+2154, the Canadian Hydrogen Intensity Mapping Experiment (CHIME) captured a fast radio burst (FRB) emission (Bochenek2020; CHIME2020b), which was precisely temporally associated with an X-ray burst (XRB; Li2021; LiXiaobo2022; Ge2022arxiv). This magnetar had just entered an active state of XRBs about half a month before FRB 20200428 occurred (Li2021; Cai2022a; Cai2022b; Mereghetti2020; Younes2020; Tavani2020; Veres2020; Ridnaia2020; Barthelmy2020; Lin2020b; Ridnaia2021; Kaneko2021), and remained active at other times as well (Israel2016MNRAS; Lin2020a; Borghese2022MNRAS; Xie2022; Xie2023arxiv; Rehan2023). Meanwhile, the dispersion measure (DM) of FRB 20200428 ($\sim 332.7$ pc cm -3) is well consistent with the distance of SGR J1935+2154 in the range of 6.612.56.6-12.5 kpc (Kothes2018; Zhou2020; Zhong2020). These discoveries indicated that magnetars could be sources of the mysterious FRB phenomena (Lorimer2007; Thornton2013).
@@
In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular, magnetars (Popov2010; Kulkarni2014; Katz2016; Connor2016; Cordes2016; Lyutikov2017), which was firstly favored by the discovery of the repeating FRB 20121102 as well as its persistent radio counterpart (Kashiyama2017; Metzger2017; Cao2017; Dai2017; Michilli2018). Then, the direct connection of FRB 20200428 with SGR J1935+2154 undoubtedly offered a smoking-gun evidence for one of the FRB/FRB-like origins, the magnetar origin, although FRB 20200428 is still distinctive due to its obviously low luminosity than its cosmological cousins (CHIME2020b; Bochenek2020).
@@
Generally, the amounts of the energy releases of the bursts and outbursts of Galactic magnetars, from both soft gamma-ray repeaters (SGRs) and anomalous X-ray pulsars (AXPs), indicate that XRBs probably correspond to sudden and local (not global) catastrophic events. Since the energy is considered to eventually come from magnetic fields, the XRBs are most likely to be caused by sudden reordering of the magnetic field structure, which either deforms the crust drastically or/and leads to a major reconfiguration of the magnetosphere. As a result, an optically thick $e^{\pm}$ fireball can be expected to form, radiating X-rays (ThompsonDuncan1995; ThompsonDuncan2001; Beloborodov2021ApJ; Lyutikov2022MNRAS). In the framework of such a starquake model, regardless of the specific X-ray radiation mechanism, some scale independent behaviors can be expected to appear in the statistics of the energy and waiting times of XRBs, just as found from earthquakes (Cheng1996). Many systems such as starquakes and earthquakes, consisting of a large number of entities that interact with each other in a complex way, can exhibit nonlinear behavior, which can be described by a self-organized criticality (SOC) theory. The complex interactions in such systems can occasionally lead to instabilities with subsequent small events. The largest events, the subsequent events, and the myriads of smaller events can finally share the same statistical property (Sornette2004book; Tsallis2009book; Aschwanden2011book).
@@
Despite a basic consensus that could appear on the XRB origin, the mechanism responsible for the FRB emission is still under great debate. In the literature, two types of models have been suggested frequently, including the curvature radiation in the magnetosphere (LuKumar2018MNRAS; YangZhang2018ApJ; YangZhang2020ApJ; LongPeer2018ApJ; WangZhang2019ApJ; KumarLu2020MNRAS) and the shock synchrotron maser radiation (Lyubarsky2014MNRAS; Beloborodov2017ApJ; Ghisellini2017MNRAS; KumarLu2017MNRAS; Lyutikov2017ApJ; Waxman2017ApJ; Metzger2019MNRAS; Plotnikov2019MNRAS; YuZou2021MNRAS). In the former case, the curvature radiation could happen in the magnetosphere because the crust fracturing forms a charge-starved region and leads to the plasma discharges. In the latter case, the shock responsible for the synchrotron maser can be driven by the collisions between different ejecta of different burst activities. In observation, although FRB 20200428 was found to be in temporal coincidence with an XRB, the deep observations implemented by the Five Hundred Meter Aperture Telescope (FAST) on some other XRBs of SGR J1935+2154 ruled out the possibility of their association with FRB emission (Lin2020c). On the one hand, this fact could just be due to the narrow beaming of FRB emission, while the XRB emission is generally isotropic. In this case, there is in principle no intrinsic difference between the different XRBs. On the other hand, reversely, the un-association with FRB emission of other XRBs may indicate that the XRB associated with FRB 20200428 is intrinsically different from the others.
@@
Figure 1: The XRBs (triangle markers) of SGR J1935+2154 detected by GBM within different observation windows (horizontal lines). The length of the lines represents the duration of the windows. The first burst in each window is labelled by a black triangle, which is dropped in our statistics since its waiting time cannot be defined. Then, the windows that only have one burst are excluded (i.e, the grey lines). The remaining windows having sufficiently large number of bursts can be basically separated into four active episodes, as labelled by different colors (see Section II.2 for active episode definition).
Figure 2: The corner plot of the shape parameter $k$ and the event rate $\lambda$ for the Weibull distribution (see Section II.1).
@@
II Temporal and energy distributions
We retrieve the burst start time, flux and fluence of SGR J1935+2154 which were detected by the Fermi/Gamma-ray Burst Monitor (GBM) from the published burst history (2014-2021, Table 4 in Lin2020a, Table 1 in Lin2020b, Table 1 in Kaneko2021, Table 1 in Rehan2023). Then, we analyze the position history of Fermi/GBM observations using POSHIST  [^1] https://heasarc.gsfc.nasa.gov/FTP/fermi/data/gbm/daily/ and calculate the uninterrupted visible time (referred to as observation window hereafter) of the GBM to the SGR J1935+2154, excluding the Earth occultation and SAA (South Atlantic Anomaly) area, using the GBM Data Tools (GbmDataTools). The start times of the observation windows and hundreds of bursts are listed in Table 1 and shown in Fig. 1. To be specific, the burst samples of SGR J1935+2154 consist of 158 observation windows, with each window having an approximate duration of an hour because GBM orbits the Earth every about 90 minutes. However some observation windows are noticeably shorter than the 90 minutes because GBM occasionally shuts down when the satellite crosses the SAA area.
II.1 The temporal clustering of XRBs
@@
where $t_{i}$ is the burst start time of the $i$-th burst. Due to the unknown waiting time for the first burst in each observation window, only 171 out of a total of 329 bursts could be analyzed (see Fig. 1). A one-sample Kolmogorov-Smirnov test is performed to examine whether the waiting times follow an exponential distribution. The test yields a distance of $D=0.28$ and $p=0.03$, indicating that a non-homogeneous model is necessary to describe the temporal behavior of the bursts. Therefore, we use the Weibull function to model the waiting times of the non-stationary Poisson process. Specifically, for an observation window of a duration $T$ which owes a number of $N$ bursts, the likelihood of this window can be calculated by (Oppermann2018),
@@
$\times\prod_{i=1}^{N-1}\mathcal{W}(t_{\mathrm{i}+1}-t_{\mathrm{i}}|k,\lambda)$
@@
Figure 3: The logarithmic waiting times, $\log(\Delta{t})$, of the XRBs v.s their occurring times, where the time of FRB 20200428 is represented by the green dashed line. The number distribution of the bursts on the waiting times are exhibited in the right panel with the histograms, while the smooth lines give the kernel density of these distributions. The histogram distributions can be approximately fitted by a Gaussian function and its mean value and variance are represented by the orange dot and the short black line in the left panel.
II.2 The cumulative distribution of the waiting times
In Fig. 3, we plot the waiting times of the XRBs of SGR J1935+2154 v.s their burst times, where an apparent clustering of bursts appear. Therefore, we can divide these bursts into 4 separated active episodes as labelled by different colors, except several bursts (grey data) outside of these episodes. Here, an active episode can be defined as a period during which more than two bursts are produced within 10 days, as previously adopted in Lin2020a; Lin2020b. The starting and ending times of the four active episodes are listed in Table 2. According to the shape of the kernel of the histogram distributions of $\log_{10}(\Delta{t})$ shown in the right panel of Fig. 3, we tentatively use a Gaussian function to fit these histogram distributions. The obtained values of the mean and variance of the Gaussians are listed in Table 2. It can be found that the mean value of the waiting times in the 3rd episode (Apr.-May 2020) during which FRB 200428 occurred is significantly shorter than those in the other episodes, indicating the particularity of the FRB episode. In addition, these near-Gaussian distributions of the waiting times make these burst activities of SGR J1935+2154 similar to the previous burst activities of magnetars (Cheng1996; Gogus1999; Gogus2000; Younes2020; Cai2022a).
@@ II.2 The cumulative distribution of the waiting times
For a more detailed description for the distribution property of the waiting times, we plot their complementary cumulative distribution (CCD) [^2] The cumulative distribution (CD) and CCD of the variable $x$ are defined in this paper as follows: $CD(x)$ $=P(\leq x)=\frac{n(\leq x)}{N}$ (5) $CCD(x)$ $=P(>x)=1-CD(x)$ where $N$ is the total number of the events. , $P(>\Delta{t})$, in the left panel of Fig. 4, where the distributions for different flux ranges are also displayed for comparison. According to the heavy-tail feature of these CCDs and following Reed2004TheDP, we try to use the double Pareto-Lognormal (dPLn) distribution to describe their general properties. The dPLn distribution, which consists of two Pareto tail distributions with a transition point in between, arises from a geometric Brownian motion process, that will be killed when the killing rate comes to constant and the state becomes equilibrium (Mitzenmacher2004ABH). Specifically, the dPLn description for the CCD of the waiting times can be expressed as (Reed2004TheDP)
@@
where $\Phi$$\Phi^{c}$ is the CD/CCD of standard normal distribution, $\nu$ and $\tau^{2}$ represent the mean and variance of the log-normal for the initial geometric Brownian motion process, and
@@
The parameters $\alpha_{1}$ and $\alpha_{2}$ determine the slope of the two Pareto tails. By using the package distribuisonsrd of Comprehensive R Archive Network (CRAN, MarçalSalvadorElena2020), we perform the dPLn fitting to the CCDs of the waiting times, as shown by the dashed line in Figure 4, and obtain $\alpha_{1}$=3.85 and $\alpha_{2}$=0.73. This indicates that, in the range of a relatively short waiting time, the CCD can be approximated by a power law with a slope of $\alpha_{2}\sim 0.73$, which could further be independent of the selection of the flux threshold and thus of the burst energy.
@@
II.3 The cumulative distribution of the fluence/flux